### v1

# MinD-3D: Reconstruct High-quality 3D objects in Human Brain

Abstract
In this paper, we introduce Recon3DMind, a groundbreaking task focused on reconstructing 3D visuals from Functional Magnetic Resonance Imaging (fMRI) signals. This represents a major step forward in cognitive neuroscience and computer vision. To support this task, we present the fMRI-Shape dataset, utilizing 360-degree view videos of 3D objects for comprehensive fMRI signal capture. Containing 55 categories of common objects from daily life, this dataset will bolster future research endeavors. We also propose MinD-3D, a novel and effective three-stage framework that decodes and reconstructs the brain’s 3D visual information from fMRI signals. This method starts by extracting and aggregating features from fMRI frames using a neuro-fusion encoder, then employs a feature bridge diffusion model to generate corresponding visual features, and ultimately recovers the 3D object through a generative transformer decoder. Our experiments demonstrate that this method effectively extracts features that are valid and highly correlated with visual regions of interest (ROIs) in fMRI signals. Notably, it not only reconstructs 3D objects with high semantic relevance and spatial similarity but also significantly deepens our understanding of the human brain’s 3D visual processing capabilities. Project page at: https://jianxgao.github.io/MinD-3D.
Figure 1: We introduce a novel task termed as Recon3DMind. Left: we showcase our data acquisition and modeling principles. Right: our proposed MinD-3D decodes fMRI signals to recover semantically matched, structurally coherent 3D objects with remarkable fidelity.
1 Introduction
Functional Magnetic Resonance Imaging (fMRI), a kind of signal that can be obtained in a non-invasive way, could capture blood changes in the human brain induced by neuronal activity. Due to its relatively easy accessibility, fMRI has been commonly used to reflect visual activities. Some recent studies [4, 5, 31, 27, 26] have successfully reconstructed high-quality images from fMRI signals either by utilizing powerful generative models e.g., diffusion models [29] or using generative transformers [3]. These approaches focus on extracting semantic features from fMRI signals, often requiring only semantic features to generate relevant high-quality images.
However, the human visual system encompasses more than the mere processing of 2D images; it possesses the remarkable ability to translate 2D projections into rich, 3D representations. This complex mechanism bestows upon us a rich perception of our world, complete with discernible aspects of size, distance, and depth. Diverging from earlier studies, our research focuses on modeling this 3D visual capability of the brain. We introduce a novel task, termed Reconstructing 3D Objects from Mind in fMRI (Recon3DMind), that leverages advanced computer vision techniques to decode and reconstruct the 3D visual information perceived by the brain from fMRI signals. This task goes beyond extracting semantic features and includes spatial and structural dimensions, which are crucial for a thorough understanding of 3D visuals.
Some literature [19, 13, 15, 20] has proved that the brain’s visual processing for 3D perception is far more complex than for 2D perception. Consequently, relying solely on semantic features is inadequate for accurately modeling the brain’s capability for 3D spatial imaging. To describe 3D objects effectively, it is essential to consider their shape and structural attributes in addition to their semantic properties. For instance, two cars of varying lengths might look identical from the front but show significant differences when viewed from the side. This observation underscores the importance of capturing the full range of spatial and structural features for a true representation of 3D objects. Therefore, our task focuses on enhancing modeling of human 3D comprehension. It calls for a sophisticated fMRI feature extractor that can discern not just semantic elements but also the spatial structure and other 3D-specific features from fMRI signals, thus facilitating a more complete and accurate recovery of 3D visual information.
Addressing the unique challenges of our Recon3DMind task, we confront a significant hurdle: the absence of a dataset pairing fMRI data with 3D visuals. To fill this void, we introduce fMRI-Shape, a pioneering dataset specifically designed for this purpose. Drawing inspiration from ZeroNVS [30], which posits that 3D objects can be effectively represented through 360-degree view videos, we adopt this methodology for fMRI data collection. In our approach, we present subjects in ShapeNet [2] with 360-degree view videos of stationary 3D objects, where the camera circumnavigates the object, completing a full rotation. This setup enables the capture of fMRI signals as subjects observe the objects from every conceivable angle. Such comprehensive visual exposure ensures the collection of detailed and precise fMRI signals, effectively capturing the entire spectrum of the object’s spatial attributes. Through intricate preprocessing steps, these recordings are transformed into multi-frame fMRI signals. The nuances and intricacies of the fMRI-Shape dataset will be elaborated further in Sec. 3.
With our specially curated fMRI-Shape dataset, we present an innovative and effective three-stage framework to extract both spatial structure and semantic features from multi-frame fMRI signals, named Mind-3D.
In the first stage, we utilize a pre-trained encoder [26] on the NSD [1] dataset to extract spatial features from the fMRI data. These features are then aggregated across multiple frames using a feature aggregation module. To ensure the biological relevance and effectiveness of these features, we align them with the visual-spatial features of the corresponding objects using contrastive learning loss. This alignment guarantees that the features we extract are not only accurate but also meaningfully connected to the brain’s visual processing.
The second stage focuses on decoding the brain’s visual activity features from the extracted fMRI data. We achieve this through a diffusion model trained within the feature space, conditioned on the fMRI features. This conditioning allows us to progressively generate accurate visual representations, effectively translating complex brain activity into comprehensible visual data.
The final stage aims to reconstruct the 3D models as they are perceived by the human brain. To accomplish this, we train a latent-adapted Argus [39] model. This model uses the visual features generated in the second stage along with the initially extracted fMRI features to create 3D objects. This process mirrors the imaging function of the brain’s visual cortex, offering a simulated insight into how the brain perceives and processes 3D objects.
To evaluate the efficacy of our model, we use both semantic and structural metrics. These metrics provide a comprehensive evaluation of the model’s ability to create structurally accurate 3D representations. Additionally, we validate the relevance of our model’s extracted features by correlating them with specific brain regions, ensuring that our model not only generates accurate representations but also does so in a way that is consistent with how the brain processes visual information.
Our contributions can be summarized as follows:
  • We for the first time define the novel task, Recon3DMind, which is meaningful for both computer vision and cognitive neuroscience communities.
  • We introduce an innovative, versatile brain decoding framework, composed of three distinct modules: a neuro-fusion encoder, a feature bridging diffusion model, and a latent adapted decoder.
  • We present fMRI-Shape, the first dataset of fMRI-3D shape pairs, to foster further research in this domain. This dataset will be accessible on our project page.
2 Related work
2.1 FMRI Decoding Methods
Current fMRI decoding methods primarily aim to reconstruct the real world in a 2D format, such as the images or videos that humans perceive. This task is quite complex, as it involves accurately extracting relevant features from fMRI signals and reconstructing a precise 2D representation. Deep learning methods, known for their impressive capabilities, are well-suited for addressing this challenge. Initial successes in this area have been demonstrated by earlier methods [18, 37, 33]. Subsequent studies [32, 7] have shown that generative models are more effective for these tasks. As a result, various diffusion models [4, 5, 31] have been employed as decoders to reconstruct visual scenes, achieving remarkable results. However, these studies have been limited to 2D visual representations. In this paper, we aim to extend the scope of fMRI decoding to 3D representations. Our goal is to directly reconstruct 3D objects from fMRI signals. To accomplish this, we propose a new framework that uses a transformer-based feature encoder for extraction and aggregation. This framework translates neural space data into visual space and employs a powerful 3D decoder to reconstruct the 3D object, utilizing features from the visual space.
2.2 Diffusion Model
Diffusion models [17, 34] are exceptional generative tools for both pixel and feature generation. As a variant, the latent diffusion model [29], equipped with an autoencoder, compresses images into lower-dimensional latent features, thereby generating a compressed version of the data rather than directly generating the data. This approach, which operates in the latent data space, significantly reduces computational requirements and enables the generation of higher-quality images with enhanced details in the latent space. In this paper, we aim to leverage the potent feature generation capabilities of diffusion models to generate visual features based on fMRI features. To achieve this, we have adapted the latent diffusion model by discarding its autoencoder structure and focusing solely on its latent component. The conditional information driving the model is derived from the fMRI features.
2.3 3D Generation
Advancements in deep generative models [36, 41] and 3D-aware image synthesis techniques [6, 38] have significantly influenced the field of 3D object generation. These models are instrumental in creating complex 3D shapes and structures, especially in architecture and virtual environment design, emphasizing discrete voxel generation and 3D modeling from 2D images. In our study, we aim to generate 3D objects from fMRI data, leveraging Argus [40], a robust 3D generative model, as our decoder and adapt it with several transformer layers. This approach facilitates an integrated method that combines fMRI and visual features produced by the previous module. The visual features serve as conditional embeddings for Argus, while the fMRI features are fused in the adapter layers. This synergistic integration aims to enhance the model’s capability to accurately reconstruct 3D objects from complex brain activities.
Figure 2: Overview of the fMRI-Shape Acquisition Process. Initially, we render each object into an 8-second long video, showcasing a 360-degree view. Following this, we capture fMRI signals in video format. The fMRI data are then processed using fMRIPrep, converting the data from 32k fs LR surface space into 2D images with dimensions of 1023 × 2514.
3 Experimental Designs and Curated Dataset
In this section, we will introduce the detailed procedure for the collection of the proposed dataset, fMRI-Shape. This dataset will be released to the community to push the research on the topic of Recon3DMind.
3.1 Subjects and Experiments
Three participants (male, aged 24-27), who were unaware of our work, were instructed to view videos depicting the gradual 360-degree rotation of 3D objects at a 60-degree pitch angle, as shown in Fig. 2. All participants had normal or corrected-to-normal visual acuity. Informed written consent was obtained from each participant, and the experimental protocol was approved by the ethical review board. To ensure the diversity of the dataset for our work, the 3D objects used in this study were derived from the ShapeNetCore [2], which includes 55 classes of objects represented as 3D point clouds, 2D images, and CAD models. We randomly selected four objects from each class in ShapeNetCore, allocating three for training and one for testing. We utilized the rendering technique from Zero123 [21] to generate an 8-second long video for each object in Blender.
For each participant, the experiment was divided into six runs. During each run, participants consecutively viewed 55 videos in a randomized order, one from each category, resulting in an approximate run duration of 8 minutes. The runs alternated between training and test data, with a 1-minute rest period after each run. Notably, each participant underwent three viewings of the testing run and a single viewing of the three training runs.
3.2 Data Acquisition and Preprocessing
MRI data were collected at Zhangjiang International Brain Imaging Centre. T1, T2 and fMRI data were acquired in a 3T scanner and a 32-channel RF head coil. T1-weighted data were scanned using MPRAGE sequence (0.8-mm isotropic resolution, TR=2500ms, TE=2.22ms, flip angle $8^{\circ}$). T2-weighted data were acquired using SPACE sequence 1(0.8-mm isotropic resolution, TR=3200ms, TE=563ms, flip angle $120^{\circ}$). Functional data were scanned using gradient-echo EPI at 2-mm isotropic resolution with whole-brain coverage (TR=800ms, TE=37ms, flip angle $52^{\circ}$, multi-band acceleration factor 8). The sampling frequency of the 3T scanner is 8Hz, so each video corresponds to a total of 10 frames of task-state fMRI signals.
Stimuli were presented using an LCD screen ($8^{\circ}\times 8^{\circ}$) positioned at the head of the scanner bed. Participants viewed the monitor via a mirror mounted on the RF coil and fixated a red central dot (0.4° × 0.4°).
Preprocessing was performed using fMRIPrep ([10, 9]). Following [27], the preprocessed functional data in 32k_fs_LR surface space were converted into 2D images and utilized for further analysis. Specifically, for each 32k_fs_LR data frame in each run, we applied z-scoring to the values across every vertex within the run. These z-scored values were then mapped onto 1023 × 2514 2D images using pycortex [11]. The ROIs for analysis were selected from the HCP-MMP atlas [12] in the 32k_fs_LR space, including “V1, V2, V3, V3A, V3B, V3CD, V4, LO1, LO2, LO3, PIT, V4t, V6, V6A, V7, V8, PH, FFC, IP0, MT, MST, FST, VVC, VMV1, VMV2, VMV3, PHA1, PHA2, PHA3”.
4 Problem Setup
Figure 3: Overview of the MinD-3D Framework. Our approach combines a Neuro-Fusion Encoder for extracting features from fMRI frames, a Feature Bridge Diffusion Model for generating visual features from these fMRI signals, and a Latent Adapted Decoder based on the Argus 3D shape generator for reconstructing 3D objects. This integrated system effectively aligns and translates brain signals into accurate 3D visual representations. Note that the CLIP encoder is only for training the model, while it is not used for the inference stage.
Recon3DMind addresses a task of paramount importance in cognitive neuroscience. Developing computational models that accurately interpret and reconstruct the human brain’s 3D comprehension can significantly advance the field of computer vision. This endeavor not only bridges a gap between cognitive neuroscience and computer vision but also promises to propel the latter field forward in unprecedented ways. In this paper, we delve into the specifics of fMRI-based 3D reconstruction, providing detailed definitions and formulas underpinning our methodology.
Consider the acquisition of a multi-frame fMRI signal denoted as $\{F\}$, where $|F|=10$. These signals correspond to both a 3D object point cloud, $\{P\}$, and a video, $\{V\}$, with $|V|=192$ frames, which the subject views. The task necessitates an efficient encoder, $E$, capable of extracting both spatial structural and semantic features from the fMRI signal. It’s crucial to note that while a single frame of the fMRI signal suffices for semantic information extraction in 2D image reconstruction, the reconstruction of 3D structures demands additional spatial structural features. Therefore, we input multiple frames of fMRI signals with the aim of extracting these comprehensive spatial structural features from the spatio-temporal data. This process is mathematically represented as: $f=E(F)$ Following the feature extraction, we employ a powerful decoder to reconstruct the original 3D point cloud $\{P\}$ based on the extracted feature $f$: $P=D(f)$ Consequently, our model can be succinctly defined as $M=\{E,D\}$, thereby operationalizing the transformation: $P=M(F)$.
In executing this model, it’s critical to fully utilize both the fMRI signal $\{F\}$ and the corresponding video $\{V\}$ to train our model $M$. To achieve this, we propose a three-stage innovative and efficient model. Each stage is meticulously designed to capture different aspects of the fMRI data and the associated visual stimuli, ensuring a comprehensive and accurate 3D reconstruction from the complex neural signals. This process not only tests the limits of current computer vision techniques but also offers valuable insights into the human brain’s processing of 3D spatial information.
5 Method
5.1 Overview
As depicted in Fig. 3, our MinD-3D model comprises three main components: 1) Neuro-Fusion Encoder: This part is responsible for extracting both semantic and structural features from the fMRI frames. 2) Feature Bridge Diffusion Model: This module generates visual features based on the features extracted by the Encoder. 3) Latent Adapted Decoder: Finally, this component recovers 3D objects, using both the fMRI features and the generated visual features.
5.2 Neuro-Fusion Encoder
Drawing from the potent Masked Image Model (MIM) [16, 26], we adopt an auto-encoder structure to extract meaningful fMRI features, trained on the UKB dataset [23]. Our encoder, termed the Neuro-Fusion Encoder (NFE), utilizes the encoder of LEA [26] as $E_{f}$ with a Feature Aggregation module $\mathcal{FA}$ to transform fMRI frames into an embedding compatible with the CLIP [28] ViT-B/32 model’s space.
Initially, we process each fMRI frame in parallel to obtain the spatial fMRI embeddings:
$\mathbf{F}_{emb}=E_{f}(\mathbf{F})$
Then, we use the Feature Aggregation module to aggregate the fMRI embeddings into fMRI latent feature $\mathbf{c}_{f}$:
$\mathbf{c}_{f}=\mathcal{FA}(\mathbf{F}_{emb})$
To align the fMRI latent feature with the vision space, we employ the ViT-B/32 CLIP vision encoder $E_{v}$ and here to encode video frames into vision embeddings. During training, we randomly select n frames as $\{\mathbf{V}\}$ and compute the average vision embeddings $V_{emb}$ as the visual latent feature $\mathbf{c}_{v}$:
$\mathbf{c}_{v}=\frac{\sum^{n}_{i=1}E_{v}(\mathbf{V_{i}})}{n}$
where n is the number of selected video frames. Then we utilize the CLIP loss to ensure the alignment between the fMRI features $\mathbf{c}_{f}$ and the visual features $\mathbf{c}_{v}$:
$\mathcal{L}_{c}=\mathcal{L}_{clip}(\mathbf{c}_{f},\mathbf{c}_{v})$
During training, we optimize the whole Neuro-Fusion Encoder with $E_{LEA}$ loading the pre-trained weights and the CLIP vision encoder frozen. It’s worth noting that the CLIP vision encoder $E_{v}$ and images $\{\mathbf{V}\}$ are just used for training, we will drop it during inference. By adopting this approach, we will get the fMRI features $\mathbf{c}_{f}$ and embedding $\mathbf{F}_{emb}$ aligned with the visual space. The fMRI features $\mathbf{c}_{f}$ will be used as the conditional information for the feature bridge diffusion model and fMRI embedding $\mathbf{F}_{emb}$ will be used in the latent adapted decoder.
Figure 4: The qualitative results generated by LEA-3D, fMRI-PTE-3D, and our method are presented. GT indicates the ground-truth 3D objects. All the objects have been rendered into a 2D format.
Methods Semantic-Level Structure-Level
[…]
LEA-3D[26] 0.787 0.371 0.353 0.527 0.562 0.575 0.624 0.372
fMRI-PTE-3D[27] 0.815 0.392 0.375 0.433 0.694 0.327 0.519 0.135
w/o Both 0.784 0.364 0.340 0.536 0.618 0.498 0.649 0.246
w/o Diffusion 0.802 0.382 0.372 0.509 0.641 0.367 0.546 0.175
w/o Contrastive 0.822 0.409 0.398 0.453 0.674 0.236 0.504 0.095
MinD-3D (full) 0.835 0.425 0.411 0.419 0.704 0.214 0.480 0.065
Table 1: Performance Comparison on the fMRI-Shape. We report average metrics for all subjects. LEA-3D and fMRI-PTE-3D represent variants of LEA and fMRI-PTE, respectively. The following 3 baselines are degenerate versions of MinD-3D.
5.3 Feature Bridge Diffusion Model
This phrase details the Neural-Visual Synthesis process using a diffusion model, named here as the Feature Bridging Diffusion Model (FBDM), following the design outlined in [29]. This module’s primary function is to facilitate the generation of the target visual latent feature $\hat{\mathbf{c}_{v}}$ from the corresponding fMRI latent feature $\mathbf{c}_{f}$.
In the forward diffusion process, we treat the visual latent feature $\mathbf{c}_{v}$ as $\mathbf{x}_{0}$ and introduce Gaussian noise into it across 100 timesteps. This approach modifies the standard 1000-timestep model to align with the settings in Mind-eye [31]:
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{t-1}+\sqrt{1-\alpha_{t}}\mathbf{\epsilon},\quad\mathbf{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$
where, $\mathbf{\epsilon}$ is the Gaussian noise. During the reverse diffusion process, the module will learn to predict and then reverse the noise introduced during the forward process. The goal is to accurately estimate the noise $\hat{\mathbf{\epsilon}}_{t}$ at each timestep, conditioned on the fMRI latent feature $\mathbf{c}_{f}$:
$\hat{\mathbf{\epsilon}}_{t}=FBDM(\mathbf{x}_{t},t,\mathbf{c}_{f})$
The fMRI latent feature $\mathbf{c}_{f}$ is mapped to the intermediate layers of the FBDM through a cross-attention operation:
CrossAttention$(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)\cdot V$, with
$Q=W_{Q}^{(i)}\varphi_{i}\left(x_{t}\right),K=W_{K}^{(i)}\tau_{\theta}(\mathbf{c}_{f}),V=W_{V}^{(i)}\tau_{\theta}(\mathbf{c}_{f}).$
Here, $\tau_{\theta}$ is the fMRI latent feature with a suitable dimension projector, $\varphi_{i}\left(x_{t}\right)$ denotes intermediate values of the Decoder and $W_{Q}^{(i)}$, $W_{K}^{(i)},W_{V}^{(i)}$ are projector matrices with learnable parameters.
So the training objective of FBDM is to minimize the difference between the actual and predicted noise:
$\mathcal{L}_{FBDM}=\mathbb{E}_{\mathbf{x},\mathbf{\epsilon},t}\left[\|\mathbf{\epsilon}-\hat{\mathbf{\epsilon}}_{t}\|^{2}\right]$
Post-training, the visual latent feature $\mathbf{c}_{v}$ will be dropped and FBDM initiates a generative phase, starting with pure noise and progressively applying the reverse diffusion process to produce the predicted visual latent feature $\hat{\mathbf{c}_{v}}$. This phase crucially maps the aligned fMRI latent features $\mathbf{c}_{f}$ to the visual space and generates the corresponding visual latent features $\hat{\mathbf{c}_{v}}$, which is used as the basic conditional information for the latent adapted decoder.
5.4 Latent Adapted Decoder
Latent Adapted Decoder is based on Argus [39], a cutting-edge 3D shape generator. It will generate the VQ latent representation of the target 3D object based on the conditional information we provide via a generative auto-regressive transformer and then get the 3D objects through the VQ decoder. Here, we design an adapter layer in the transformer block to modify the module using the fMRI embeddings $\mathbf{F}_{emb}$.
Concretely, to directly incorporate the fMRI feature into the 3D generation process, we adapt the Multihead Self-Attention (MSA) in the standard transformer block into a double attention layer. This layer consists of the original MSA and an adapter layer. The adapter layer calculates multi-head cross-attention using the intermediate values of the module. We then obtain the adapted result using a learnable parameter $\lambda$ as the coefficient:
$x=\lambda\cdot\textsc{CA}(Q,K_{f},V_{f})+(1-\lambda)\cdot\textsc{SA}(Q,K,V)$
where CA and SA indicate cross-attention and self-attention separately; $Q,K,V$ denotes the representation of 3d latent features with a Linear projection; $K_{f},V_{f}$ are the representation of fMRI embeddings with a linear projection. The adapter layer is added every T transformer block.
Our decoder computes the VQ latent vector for the target object:
$p(\mathbf{z})=\prod_{i=1}^{m}p\left(\mathbf{z}_{i}\mid\hat{\mathbf{c}_{v}},\mathbf{F}_{emb},\mathbf{z}_{<i}\right)$
Finally, the 3D VQ-decoder decodes this latent vector $\mathbf{z}$ to reconstruct the final point clouds $\mathbf{P}$:
$\{\mathbf{P}\}=D(\mathbf{z})$
The decoder’s optimization objective is defined by the negative log-likelihood loss:
$\mathcal{L}_{nll}=\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}\left[-\log p(\mathbf{z})\right]$
During training, we train the module separately using pre-training weights, after training the neuro-fusion encoder and the feature bridge diffusion model.
6 Expriment
6.1 Metrics
In order to evaluate the performance of the new task and our model comprehensively, we adopt the following metrics in two aspects, semantic-level and structure-level:
Semantic Level. To evaluate the semantic quality of our model, we adopt standard metrics used in previous 2D fMRI literature [24, 4, 5, 25, 22], namely pixel-level correlation (PixCorr) and N-way top-K accuracy. These two metrics are calculated by comparing the reconstructed images with the ground truth (GT) images. In this process, we render both the reconstructed objects and the GT objects into images separately, calculate the metrics for each frame, and then derive the average value as the final metric numbers.
Structure Level. Besides evaluating the semantic quality of our model, it is also essential to assess its performance in terms of spatial structure, which we do in both 3D and 2D. In 3D, we utilize common metrics used in 3D reconstruction: Chamfer Distance (CD), Hausdorff Distance (HD), and Earth Mover’s Distance (EMD). For the 2D evaluation, we render the 3D objects onto a 2D plane and compute the frame-averaged pixel-level correlation, Learned Perceptual Image Patch Similarity (LPIPS), and Structural Similarity (SSIM).
6.2 Implementation Details.
As described in Section 3, each data pair in the fMRI-Shape dataset consists of 10 fMRI frames and the corresponding 192 images. To maximize the utilization of this dataset, we randomly select 6 fMRI frames for training, which also serves as a form of data augmentation. These 6 frames, initially 1023 × 2514 2D fMRI images, are resized to 256 × 256. For the image data, we set the number of selected video frames (n) to 4, meaning we randomly choose 4 frames for subsequent training, and each of these images is resized to 224 × 224. During inference, we select only the middle 6 fMRI frames and do not input any additional images. In all experiments, we train the model using data from all 3 subjects and then calculate the metrics for each subject.
Regarding the model structure, our neuro-fusion encoder comprises 24 transformer layers with 1024 feature dimensions and 16 self-attention heads. Utilizing the CLIP ViT-B/32 model, our Feature Aggregation module maps features to a 512-dimensional space. Consequently, the feature bridge diffusion model is configured with a feature dimension of 512 for cross-attention and 8 self-attention heads, with the number of diffusion channels set to 224. The latent adapted decoder is designed with 32 transformer layers, 3072 feature dimensions, and 16 self-attention heads. We set T=8, which means that we add an adapter layer every 8 transformer blocks.
In terms of model training, the first stage involves optimizing both the neuro-fusion encoder and the feature bridge diffusion model. We initialize the neuro-fusion encoder with pre-trained weights, and the feature bridge diffusion model is trained from scratch. In the second stage, we fix the parameters of the first two stages and focus on optimizing the latent adapted decoder. We use the AdamW optimizer with a learning rate of 5e-5 for all experiments and train the model with a batch size of 128 for all subjects.
6.3 Experimental Results
As the first work to model 3D imaging in the human brain, we compare our comprehensive model, MinD-3D, with several baselines, including adaptations of LEA and fMRI-PTE using the same 3D decoder as in MinD-3D. Tab. 1 presents the averaged metrics at both the Structural and Semantic Levels for all subjects. MinD-3D demonstrates superior performance across all metrics: achieving approximately 83.5% accuracy in 2-way-top-1, 42.5% accuracy in 10-way-top-1, a pixel-level correlation of 0.411, and 0.419 in LPIPS. These results indicate that our model not only generates objects with high semantic accuracy but also excels in structural similarity, outperforming baseline methods.
Qualitative results from LEA-3D, fMRI-PTE-3D, and our method are showcased in Fig. 4. MinD-3D consistently produces 3D objects that are structurally very similar to their real counterparts, while maintaining semantic integrity in most instances. A notable example is the follower vase case, where despite the complexity of its semantics and structure, our model effectively captures and reproduces key features. This underscores the robustness of our model in handling this novel task and its capability to produce faithful realizations.
6.4 Analytical and Ablation Study
We further conduct some analytical and ablation studies to evaluate the effectiveness of our model.
Figure 5: Quality of the extracted features. We map the features to brain regions via ridge regression modeling, with voxel-wise Pearson’s correlation coefficients computed between them. The visual function ROI is highlighted in red, indicating correlations in the range of 0 to 0.8. In contrast, other regions are depicted in blue, representing correlations within a range of 0 to 0.5.
Analysis of feature extracted. To understand our feature extractor from a biological perspective, following [14, 37, 35, 24], we employ a linear encoding model here to project the extracted fMRI feature $c_{f}$ to brain activity. For this purpose, we select subject 1 as our representative. Specifically, we employ ridge regression to learn the mapping relationship between $c_{f}$ and fMRI signals in the training set. We then map $c_{f}$ extracted from the test set to the fMRI space and calculate the Pearson correlation coefficient. As illustrated in Fig. 5, we use red to indicate the positive correlation between visual ROIs. For other regional positive correlations, we use blue. A high positive correlation is evident between $c_{f}$ and visual ROIs, indicating that we have successfully extracted valid features. Additionally, we conducted a similar analysis with the visual latent feature $c_{v}$ and visualized its positive correlations. Notably, the correlation of $c_{f}$ is superior to that of $c_{v}$. Furthermore, we calculated the correlation between $c_{f}$ and $c_{v}$, which is very high. This high correlation suggests that our comparative learning is effective and that the fMRI feature space and image space are well-aligned.
Ablation study of fMRI frames. During training, we randomly select 6 fMRI frames, denoted as $F_{n}$, as the input to our model. To explore the impact of different $F_{n}$ values on our model’s performance, we conducted an assessment across various $F_{n}$ values. We report and compare average results for metrics such as 2-way-top-1, 10-way-top-1, SSIM, and CD, focusing on the performance across three subjects. As shown in Tab. 2, $F_{n}=6$ emerges as the optimal choice. This operation not only acts as data augmentation but also, due to the inherent noise in fMRI signals, choosing the middle portion for inference helps in reducing the introduction of noise.
7 Conclusion
$F_{n}$ Semantic-level Structure-Leve
[…]
1 0.674 0.291 0.472 0.853
3 0.762 0.365 0.604 0.526
6 0.835 0.425 0.704 0.214
8 0.829 0.401 0.702 0.225
10 0.818 0.379 0.684 0.273
Table 2: Ablation results for the fMRI frames. It indicates that $F_{n}=6$ is the best choice for MinD-3D.
In this paper, for the first time, we introduce the task of 3D reconstruction based on fMRI signals, accompanied by a dedicated dataset. Technically, we propose a novel three-stage framework tailored for this groundbreaking task. In the initial stage, our model adeptly extracts features from fMRI frames. Subsequently, in the second stage, it employs a diffusion module to translate these fMRI features into the image domain. The final stage involves the transformation of these image features into 3D models using an advanced 3D generation model. Our experimental results and analyses demonstrate the efficacy of our model in accurately extracting fMRI features and converting them into their corresponding 3D objects. This work not only establishes a new domain in neuroimaging and 3D reconstruction but also opens avenues for further research in understanding and visualizing neural representations in 3D vision.
Appendix A More Implementation Details
To more effectively demonstrate the details of our model and its implementation, this section introduces the necessary preliminary knowledge and provides a detailed description of our neuro-fusion encoder and latent adapted decoder.
A.1 Neuro-Fusion Encoder
Our encoder consists of the encoder from the pre-trained LEA [26] and a feature aggregation module. LEA, a transformer-based auto-encoder, encodes fMRI signals and reconstructs them as depicted in Fig. A.1. Unlike the original Masked Autoencoder (MAE) [16], which only masks some patch-based latent codes, LEA masks all such codes. This enables the decoder to learn to reconstruct fMRI signals from the learned [CLS] token during training, thereby facilitating the development of a dense fMRI representation within the [CLS] token.
LEA is trained on the UKB dataset [23]. As shown in Fig. A.1, when LEA is directly applied to our fMRI-Shape dataset, it yields rough reconstructions. However, these reconstructions are not highly precise, primarily due to variations in scanners, subjects, and tasks.
Figure A.1: Architecture of LEA. During training, LEA masks all the patch-based latent codes and enables the decoder to learn to reconstruct fMRI signals from the learned [CLS] token.
Given that our input fMRI signals are in video form, to efficiently extract both semantic and structural features from this complex data, we opt to fine-tune the encoder with a small learning rate during the initial stage of training with hyperparameters provided in Tab. A.1.
parameter value parameter value
patch size 16 encoder depth 24
embedding dim 1024 encoder heads 16
mask ratio 1.0 fMRI frames 6
mlp ratio 0.99 image frames 4
Table A.1: Hyperparameters used in neuro-fusion encoder.
A.2 Latent Adapted Decoder
Our decoder is adapted from Argus [39]. The original Argus model enhances the quality and diversity of generated 3D shapes by incorporating tri-plane features as latent representations and a discrete codebook for efficient quantization, leveraging the power of transformers for multi-modal conditional generation. It generates latent codes in an auto-regressive manner, similar to VQGAN [8], based on conditional features. Subsequently, it decodes the quantized vectors into point clouds.
In our latent adapted decoder, we add an adapter layer after every 8 transformer blocks to fuse the fMRI feature. During the second stage of training, we focus on optimizing the parameters in the adapter layers and the learnable parameter $\lambda$. The hyperparameters of our latent adapted decoder are detailed in Tab. A.2.
parameter value parameter value
codebook size 8192 encoder depth 32
codebook dim 512 encoder heads 16
embedding dim 3072 adapter layers 4
mlp ratio 0.99 sequence length 1024
Table A.2: Hyperparameters used in latent adapted decoder.
Appendix B FMRI-Shape Details
To better illustrate our proposed dataset, fMRI-Shape, we provide a detailed description in this section.
Figure B.1: Examples of 360-Degree View Videos. These videos effectively showcase the intricate details, lighting, and shadows of the 3D objects. Presented in a sequence from left to right, they depict the order in which the subjects viewed them. Each video boasts a high resolution of 1024×1024, providing a clear and detailed visual experience.
B.1 More Experimental Details
To clarify the experimental details, we showcase our specially rendered 360-degree view videos. In Fig. B.1, six frames are selected from four videos. The camera in each video encircles the object, completing a full 360-degree clockwise rotation, forming an 8-second video with 192 frames. These high-resolution videos (1024×1024 pixels) vividly capture the intricate details, lighting effects, and shadows of the 3D objects. The arrangement of the videos mirrors the order in which they were viewed by the subjects. Displayed on an LCD screen, these videos enable the subjects to observe and comprehend all facets of the 3D objects. This viewing strategy is crucial, as it aids the subjects’ brains in accurately perceiving the objects, a fundamental aspect for gathering precise and high-quality fMRI signals that correspond to the visual stimuli observed by the subjects.
B.2 More Preprocessing Details
In preparing our data for training, we provide a visualization that clearly shows how we identify the visual Regions of Interest (ROIs) in our data. The preprocessing is applied to all fMRI frames.
Figure B.2: Selection Process of Visual ROIs: On the left, the fMRI is depicted in its original signal form, while on the right, the corresponding area is transformed into a 2D image, demonstrating the conversion process.
Refer to Fig. B.2. The first row displays the raw fMRI signals and their 2D image formation. In the second row, on the left, you can see the appearance of the fMRI signal and the locations of the ROIs. On the right, there is a 2D image illustrating where these ROIs are located within the signal, which has been converted from a format known as 32k_fs_LR.
After identifying these ROIs in the 1023×2514 2D fMRI data, we resize them to 256×256, as shown in Fig. B.3. It’s important to note that our Mind-3D model uses these resized, single-channel 2D fMRI signals as its input.
Figure B.3: Converting ROIs in fMRI Signal
Appendix C More Results
Subject Semantic-Level Structure-Level
[…]
sub-001 0.842 0.424 0.411 0.421 0.704 0.212 0.478 0.069
sub-002 0.836 0.424 0.412 0.420 0.705 0.218 0.483 0.062
sub-003 0.829 0.427 0.410 0.416 0.703 0.212 0.479 0.064
Table C.1: Quantitative results for all subjects.
Qualitative results from different views. In Fig. C.1, we showcase a collection of 2D rendered images of 3D objects by MinD-3D from various viewpoints for subject-1 in fMRI-Shape, closely resembling the ground truth (GT) objects both in semantics and structure.
Quantitative results for all subjects. Our proposed dataset, fmri-shape, includes data from three subjects. To demonstrate the robustness of our model, we present comprehensive quantification results for all three subjects. Tab. C.1 showcases these full metrics, indicating that our model consistently performs well across different subjects. These results underscore the model’s robustness in handling subject variability.
Figure C.1: Qualitative results for subject-1 from various viewpoints.
Appendix D Limitations and Future Works
As the first attempt at the task of Reconstructing 3D Objects from Mind in fMRI (Recon3DMind), our MinD-3D has successfully reconstructed 3D objects that are semantically and structurally similar to the originals. However, the reconstructed 3d objects lack detailed texture, color, and other aspects compared to the original point clouds. To overcome this limitation, our future work will focus on developing methods to more effectively extract texture and color details from 3D objects within fMRI signals. We aim to refine our reconstruction method to create more accurate and detailed replications of the original 3D objects, thereby enhancing the fidelity of the reconstructed models and capturing the complexity and richness of the original designs. Regarding our dataset, we recognize that relying solely on the fMRI-Shape dataset is insufficient for the task’s full scope. A notable challenge is mitigating dataset noise due to the cognitive biases individuals might hold towards certain objects. Therefore, it’s necessary for us to expand our fMRI-Shape dataset by including more 3D objects and collecting data from additional subjects, with the hope of advancing research in our proposed Recon3DMind field.

### v3

@@
 Abstract
-In this paper, we introduce Recon3DMind, a groundbreaking task focused on reconstructing 3D visuals from Functional Magnetic Resonance Imaging (fMRI) signals.
-This represents a major step forward in cognitive neuroscience and computer vision.
-To support this task, we present the fMRI-Shape dataset, utilizing 360-degree view videos of 3D objects for comprehensive fMRI signal capture.
-Containing 55 categories of common objects from daily life, this dataset will bolster future research endeavors.
-We also propose MinD-3D, a novel and effective three-stage framework that decodes and reconstructs the brain’s 3D visual information from fMRI signals.
-This method starts by extracting and aggregating features from fMRI frames using a neuro-fusion encoder, then employs a feature bridge diffusion model to generate corresponding visual features, and ultimately recovers the 3D object through a generative transformer decoder.
-Our experiments demonstrate that this method effectively extracts features that are valid and highly correlated with visual regions of interest (ROIs) in fMRI signals.
-Notably, it not only reconstructs 3D objects with high semantic relevance and spatial similarity but also significantly deepens our understanding of the human brain’s 3D visual processing capabilities.
+In this paper, we introduce Recon3DMind, an innovative task aimed at reconstructing 3D visuals from Functional Magnetic Resonance Imaging (fMRI) signals, marking a significant advancement in the fields of cognitive neuroscience and computer vision.
+To support this pioneering task, we present the fMRI-Shape dataset, which includes data from 14 participants and features 360-degree videos of 3D objects to enable comprehensive fMRI signal capture across various settings, thereby laying a foundation for future research.
+Furthermore, we propose MinD-3D, a novel and effective three-stage framework specifically designed to decode the brain’s 3D visual information from fMRI signals, demonstrating the feasibility of this challenging task.
+The framework begins by extracting and aggregating features from fMRI frames through a neuro-fusion encoder, subsequently employs a feature bridge diffusion model to generate visual features, and ultimately recovers the 3D object via a generative transformer decoder.
+We assess the performance of MinD-3D using a suite of semantic and structural metrics and analyze the correlation between the features extracted by our model and the visual regions of interest (ROIs) in fMRI signals.
+Our findings indicate that MinD-3D not only reconstructs 3D objects with high semantic relevance and spatial similarity but also significantly enhances our understanding of the human brain’s capabilities in processing 3D visual information.
 Project page at: https://jianxgao.github.io/MinD-3D.
-Figure 1: We introduce a novel task termed as Recon3DMind.
-Left: we showcase our data acquisition and modeling principles.
-Right: our proposed MinD-3D decodes fMRI signals to recover semantically matched, structurally coherent 3D objects with remarkable fidelity.
+Keywords:
+fMRI 3D vision diffusion model
 1 Introduction
@@
 Due to its relatively easy accessibility, fMRI has been commonly used to reflect visual activities.
-Some recent studies [4, 5, 31, 27, 26] have successfully reconstructed high-quality images from fMRI signals either by utilizing powerful generative models e.g., diffusion models [29] or using generative transformers [3].
+Some recent studies [6, 7, 41, 36, 34] have successfully reconstructed high-quality images from fMRI signals by utilizing powerful generative models [39, 3].
 These approaches focus on extracting semantic features from fMRI signals, often requiring only semantic features to generate relevant high-quality images.
+Figure 1: Overview of Recon3DMind task, showcasing the fMRI-Shape dataset collection process with 14 participants observing 360-degree view videos of 3D objects, and MinD-3D framework for reconstructing 3D objects from fMRI signals.
 However, the human visual system encompasses more than the mere processing of 2D images; it possesses the remarkable ability to translate 2D projections into rich, 3D representations.
@@
 This task goes beyond extracting semantic features and includes spatial and structural dimensions, which are crucial for a thorough understanding of 3D visuals.
-Some literature [19, 13, 15, 20] has proved that the brain’s visual processing for 3D perception is far more complex than for 2D perception.
-Consequently, relying solely on semantic features is inadequate for accurately modeling the brain’s capability for 3D spatial imaging.
-To describe 3D objects effectively, it is essential to consider their shape and structural attributes in addition to their semantic properties.
-For instance, two cars of varying lengths might look identical from the front but show significant differences when viewed from the side.
-This observation underscores the importance of capturing the full range of spatial and structural features for a true representation of 3D objects.
-Therefore, our task focuses on enhancing modeling of human 3D comprehension.
-It calls for a sophisticated fMRI feature extractor that can discern not just semantic elements but also the spatial structure and other 3D-specific features from fMRI signals, thus facilitating a more complete and accurate recovery of 3D visual information.
-Addressing the unique challenges of our Recon3DMind task, we confront a significant hurdle: the absence of a dataset pairing fMRI data with 3D visuals.
-To fill this void, we introduce fMRI-Shape, a pioneering dataset specifically designed for this purpose.
-Drawing inspiration from ZeroNVS [30], which posits that 3D objects can be effectively represented through 360-degree view videos, we adopt this methodology for fMRI data collection.
-In our approach, we present subjects in ShapeNet [2] with 360-degree view videos of stationary 3D objects, where the camera circumnavigates the object, completing a full rotation.
-This setup enables the capture of fMRI signals as subjects observe the objects from every conceivable angle.
-Such comprehensive visual exposure ensures the collection of detailed and precise fMRI signals, effectively capturing the entire spectrum of the object’s spatial attributes.
-Through intricate preprocessing steps, these recordings are transformed into multi-frame fMRI signals.
-The nuances and intricacies of the fMRI-Shape dataset will be elaborated further in Sec.
+Some literature [15, 16, 25] has shown that the brain’s visual processing mechanisms for 3D perception are significantly more complex than those for 2D perception.
+This increased complexity is manifested in the differential activation of brain regions during 3D visualization tasks [14, 23].
+Consequently, a reliance solely on semantic features falls short of accurately modeling the brain’s capacity for 3D spatial imaging.
+Effective description of 3D objects necessitates the consideration of their shape and structural attributes, in addition to semantic properties.
+For example, two cars may appear identical when viewed from the front, yet present substantial differences in length when observed from the side.
+This highlights the critical need to capture the full range of spatial and structural features for an authentic representation of 3D objects.
+Accordingly, our work aims to advance the modeling of human 3D comprehension.
+It necessitates the development of an advanced fMRI feature extractor capable of discerning not only semantic elements but also spatial structures and other 3D-specific characteristics from fMRI signals.
+This approach aims at enabling a more complete and accurate reconstruction of 3D visual information.
+Addressing the unique challenges of our Recon3DMind task, we confront a significant hurdle: the absence of a dataset that pairs fMRI data with 3D visuals.
+To fill this void, we introduce fMRI-Shape, a pioneering resource specifically designed for this purpose, featuring data from 14 participants and supporting various experimental settings.
+Inspired by ZeroNVS [40], which suggests that 3D objects can be effectively represented through 360-degree view videos, we adopt a similar methodology for our fMRI data collection.
+In our approach, participants are presented with 360-degree view videos of stationary 3D objects from ShapeNet [2].
+These videos feature the camera orbiting the object, completing a full rotation, thereby providing subjects with a comprehensive view of the objects from all possible angles.
+This method ensures the capture of detailed and precise fMRI signals as subjects observe the objects, effectively encapsulating the full range of the object’s spatial attributes.
+Through meticulous preprocessing, these recordings are converted into multi-frame fMRI signals, capturing a rich dataset that allows for a nuanced analysis.
+The complexities and specific details of the fMRI-Shape dataset will be further elaborated in Sec.
 3.
@@
 The second stage focuses on decoding the brain’s visual activity features from the extracted fMRI data.
-We achieve this through a diffusion model trained within the feature space, conditioned on the fMRI features.
+We achieve this through a transformer-based diffusion model trained within the feature space, conditioned on the fMRI features.
 This conditioning allows us to progressively generate accurate visual representations, effectively translating complex brain activity into comprehensible visual data.
@@
 This process mirrors the imaging function of the brain’s visual cortex, offering a simulated insight into how the brain perceives and processes 3D objects.
-To evaluate the efficacy of our model, we use both semantic and structural metrics.
-These metrics provide a comprehensive evaluation of the model’s ability to create structurally accurate 3D representations.
-Additionally, we validate the relevance of our model’s extracted features by correlating them with specific brain regions, ensuring that our model not only generates accurate representations but also does so in a way that is consistent with how the brain processes visual information.
+To evaluate the efficacy of our model, we employ various semantic and structural metrics.
+These metrics offer a comprehensive assessment of the model’s ability to generate structurally accurate 3D representations.
+Additionally, we validate the relevance of the features extracted by our model by correlating them with specific brain regions.
+This ensures that our model not only produces accurate representations but also does so in a manner consistent with the brain’s processing of visual information.
 Our contributions can be summarized as follows:
-  • We for the first time define the novel task, Recon3DMind, which is meaningful for both computer vision and cognitive neuroscience communities.
-  • We introduce an innovative, versatile brain decoding framework, composed of three distinct modules: a neuro-fusion encoder, a feature bridging diffusion model, and a latent adapted decoder.
-  • We present fMRI-Shape, the first dataset of fMRI-3D shape pairs, to foster further research in this domain. This dataset will be accessible on our project page.
+  • For the first time, we propose and demonstrate the feasibility of the novel task, Recon3DMind, which holds significance for both the computer vision and cognitive neuroscience communities.
+  • We present fMRI-Shape, the first dataset of fMRI-3D shape pairs, across various settings, to foster further research in this domain. This dataset will be accessible on our project page.
+  • We introduce an innovative, versatile brain decoding framework MinD-3D, composed of three distinct modules: a neuro-fusion encoder, a feature-bridging diffusion model, and a latent-adapted decoder.
 2 Related work
 2.1 FMRI Decoding Methods
-Current fMRI decoding methods primarily aim to reconstruct the real world in a 2D format, such as the images or videos that humans perceive.
-This task is quite complex, as it involves accurately extracting relevant features from fMRI signals and reconstructing a precise 2D representation.
-Deep learning methods, known for their impressive capabilities, are well-suited for addressing this challenge.
+Current fMRI decoding methods primarily focus on reconstructing the vision perception in a 2D format, such as the images or videos perceived by humans.
+This task is complex, as it requires accurately extracting relevant features from fMRI signals and reconstructing a precise 2D representation.
+Deep learning methods, known for their impressive capabilities, are well-suited to address this challenge.
 Initial successes in this area have been demonstrated by earlier methods [19, 51, 43].
-Subsequent studies [32, 7] have shown that generative models are more effective for these tasks.
-As a result, various diffusion models [4, 5, 31] have been employed as decoders to reconstruct visual scenes, achieving remarkable results.
-However, these studies have been limited to 2D visual representations.
-In this paper, we aim to extend the scope of fMRI decoding to 3D representations.
+Subsequent studies [42, 10] have shown that generative models are particularly effective for these tasks, leading to the employment of various diffusion models [6, 7, 41, 45] as decoders to reconstruct visual scenes, achieving remarkable results.
+However, these studies have been limited to 2D visual representations and the related vision ROIs.
+In this paper, we aim to extend the scope of fMRI decoding to 3D representations, involving more vision ROIs.
 Our goal is to directly reconstruct 3D objects from fMRI signals.
-To accomplish this, we propose a new framework that uses a transformer-based feature encoder for extraction and aggregation.
-This framework translates neural space data into visual space and employs a powerful 3D decoder to reconstruct the 3D object, utilizing features from the visual space.
-2.2 Diffusion Model
+To accomplish this, we propose a new framework that employs a transformer-based feature encoder for extraction and aggregation.
+This framework translates neural space data into visual space and utilizes a powerful 3D decoder to reconstruct the 3D object, leveraging features from the visual space.
+2.2 Diffusion Models
 Diffusion models [18, 44] are exceptional generative tools for both pixel and feature generation.
-As a variant, the latent diffusion model [29], equipped with an autoencoder, compresses images into lower-dimensional latent features, thereby generating a compressed version of the data rather than directly generating the data.
-This approach, which operates in the latent data space, significantly reduces computational requirements and enables the generation of higher-quality images with enhanced details in the latent space.
-In this paper, we aim to leverage the potent feature generation capabilities of diffusion models to generate visual features based on fMRI features.
-To achieve this, we have adapted the latent diffusion model by discarding its autoencoder structure and focusing solely on its latent component.
+As a variant, the latent diffusion model [39], equipped with an autoencoder, compresses images into lower-dimensional latent features, thereby generating a compressed version of the data rather than directly generating the data itself.
+Dit [33] replaces the backbone of diffusion models with transformers, which will improve the performance and scalability of these models.
+This approach, operating in the latent space, significantly reduces computational requirements and enables the generation of higher-quality images with enhanced details in the latent space.
+In this paper, we aim to leverage the potent feature-generation capabilities of diffusion models to generate visual features based on fMRI features.
+To achieve this, we have adapted a transformer-based diffusion model, focusing solely on its latent component.
 The conditional information driving the model is derived from the fMRI features.
 2.3 3D Generation
-Advancements in deep generative models [36, 41] and 3D-aware image synthesis techniques [6, 38] have significantly influenced the field of 3D object generation.
-These models are instrumental in creating complex 3D shapes and structures, especially in architecture and virtual environment design, emphasizing discrete voxel generation and 3D modeling from 2D images.
-In our study, we aim to generate 3D objects from fMRI data, leveraging Argus [40], a robust 3D generative model, as our decoder and adapt it with several transformer layers.
-This approach facilitates an integrated method that combines fMRI and visual features produced by the previous module.
-The visual features serve as conditional embeddings for Argus, while the fMRI features are fused in the adapter layers.
-This synergistic integration aims to enhance the model’s capability to accurately reconstruct 3D objects from complex brain activities.
+3D generation can be accomplished through various methods [35, 56, 48, 26].
+Some methods [49, 48] employ 3D Gaussian splatting [24] for this purpose.
+Other studies [50, 54] utilize diffusion models to generate multi-view representations of objects, subsequently constructing 3D models.
+Additionally, traditional and direct approaches leverage autoregressive methods [8, 21, 35] for 3D object generation.
+In our study, we aim to generate 3D objects from fMRI data, using Argus [35], a robust 3D generative model, as our decoder, which we enhance with several transformer layers.
+This approach integrates fMRI and visual features produced by the preceding module.
+The visual features act as conditional embeddings for Argus, while the fMRI features are fused within the adapter layers.
+This synergistic integration seeks to augment the model’s capability to accurately reconstruct 3D objects from complex brain activities.
+Figure 2: Comparing fMRI-Shape with other 2D fMRI datasets.
+As the first 3D fMRI dataset, fMRI-Shape features a larger number of participants and frames, providing ample support for experiments in our proposed novel task and further research.
+Table 1: Details of fMRI-Shape Dataset.
+It ensures a balanced representation of male and female participants.
+Specifically, the Across-Person Set (AP Set) includes fMRI data sampled from different participants for Core test data.
+The Across-Person & Across-Class Set (APAC Set) involves different participants viewing new 3D objects.
+These two subsets facilitate evaluations for model generalization.
+Participant Males/Females Category Objects Frames
+Core Set 8 4/4 13 1404 14040
+AP Set 2 1/1 13 104 1040
+APAC Set 4 2/2 55 220 2200
+Total 14 7/7 55 1624 123200
+3 Experimental Designs and Curated Dataset
+In this section, we will introduce the detailed procedure for collecting the proposed dataset, fMRI-Shape.
+The scale of fMRI-Shape compared to other datasets (NSD [1], BOLD5000 [4], GOD [19], Video-fMRI [51]) is illustrated in Fig.
+2, while the specific details of fMRI-Shape are provided in Tab.
+1.
+This dataset are released to the community to push the research on the topic of Recon3DMind.
 Figure 3: Overview of the fMRI-Shape Acquisition Process.
 Initially, we render each object into an 8-second long video, showcasing a 360-degree view.
-Following this, we capture fMRI signals in video format.
-The fMRI data are then processed using fMRIPrep, converting the data from 32k fs LR surface space into 2D images with dimensions of 1023 × 2514.
-3 Experimental Designs and Curated Dataset
-In this section, we will introduce the detailed procedure for the collection of the proposed dataset, fMRI-Shape.
-This dataset will be released to the community to push the research on the topic of Recon3DMind.
+Subsequent fMRI signal capture is performed in video format, followed by data processing with fMRIPrep to convert signals from 32k_fs_LR surface space into 2D images of dimensions 1023 $\times$ 2514.
+Individual differences observed in the dataset, as highlighted in the middle part, underscore the challenges in generalizing these findings.
+On the rights, regions of interest (ROIs) are transformed into 256 $\times$ 256 image.
 3.1 Subjects and Experiments
-Three participants (male, aged 24-27), who were unaware of our work, were instructed to view videos depicting the gradual 360-degree rotation of 3D objects at a 60-degree pitch angle, as shown in Fig.
-2.
-All participants had normal or corrected-to-normal visual acuity.
-Informed written consent was obtained from each participant, and the experimental protocol was approved by the ethical review board.
-To ensure the diversity of the dataset for our work, the 3D objects used in this study were derived from the ShapeNetCore [2], which includes 55 classes of objects represented as 3D point clouds, 2D images, and CAD models.
-We randomly selected four objects from each class in ShapeNetCore, allocating three for training and one for testing.
-We utilized the rendering technique from Zero123 [21] to generate an 8-second long video for each object in Blender.
-For each participant, the experiment was divided into six runs.
-During each run, participants consecutively viewed 55 videos in a randomized order, one from each category, resulting in an approximate run duration of 8 minutes.
-The runs alternated between training and test data, with a 1-minute rest period after each run.
-Notably, each participant underwent three viewings of the testing run and a single viewing of the three training runs.
+FMRI-Shape encompasses data from 14 participants unfamiliar with the objectives of our study.
+All participants possessed normal or corrected-to-normal visual acuity.
+Informed written consent was obtained from each participant, and the experimental protocol received approval from the ethical review board.
+To ensure the diversity of the dataset, the 3D objects used in this study were derived from the ShapeNetCore [2], which includes 55 classes of objects.
+We utilized the rendering technique from Zero123 [27] to generate an 8-second video for each object using Blender.
+These videos illustrate the gradual 360-degree rotation of 3D objects at a 60-degree pitch angle, as exemplified in Fig.
+3.
+Core The core set of fMRI-Shape includes data from 8 participants: 4 males and 4 females, aged 21 to 29 (Participants No.
+1-8).
+We selected a total of 1404 objects across 13 common categories frequently employed in various 3D reconstruction literature [5, 46, 22] within ShapeNetCore.
+For each category, 100 objects were allocated for training and 8 for testing, totaling 108 objects per category.
+The experiment lasted approximately 4.5 hours per participant, divided into 32 sessions.
+In each session, participants viewed 48 videos in a randomized order, interspersed with 2.4 second rest periods, covering one object from each category, resulting in a session duration of around 9 minutes.
+Sessions alternated between training and test data, with a 1-hour break after every 8 sessions to prevent fatigue and maintain data quality.
+Notably, the testing objects were viewed twice by each participant.
+Across-Person Set (AP Set) The AP Set, designed for Out-Of-Distribution (OOD) testing, includes fMRI data from 2 participants (1 male aged 24 and 1 female aged 26, Participants No.
+9 and 10), sampled while they were viewing the test objects from the Core set.
+Across-Person & Across-Class Set (APAC Set) The APAC Set represents a more challenging component than the AP Set.
+It includes data from 4 participants: 2 males and 2 females, aged 22 to 26 (Participants No.
+11-14).
+We randomly selected 4 objects from each of the 55 categories in ShapeNetCore, distinct from those in the Core, for a total of 220 objects.
+Each participant was exposed to these 220 objects across 6 sessions.
+As shown in the middle part of Fig.
+3, we can observe individual differences, which will pose significant challenges for generalization.
+Thus, the AP and APAC sets are crucial for OOD testing and will serve as important benchmarks to assess the generalization capabilities of the 3D decoding model.
 3.2 Data Acquisition and Preprocessing
-MRI data were collected at Zhangjiang International Brain Imaging Centre.
-T1, T2 and fMRI data were acquired in a 3T scanner and a 32-channel RF head coil.
+T1 and fMRI data were acquired in a 3T scanner and a 32-channel RF head coil.
 T1-weighted data were scanned using MPRAGE sequence (0.8-mm isotropic resolution, TR=2500ms, TE=2.22ms, flip angle $8^{\circ}$).
-T2-weighted data were acquired using SPACE sequence 1(0.8-mm isotropic resolution, TR=3200ms, TE=563ms, flip angle $120^{\circ}$).
 Functional data were scanned using gradient-echo EPI at 2-mm isotropic resolution with whole-brain coverage (TR=800ms, TE=37ms, flip angle $52^{\circ}$, multi-band acceleration factor 8).
-The sampling frequency of the 3T scanner is 8Hz, so each video corresponds to a total of 10 frames of task-state fMRI signals.
+The sampling frequency of the 3T scanner is 1.25Hz, so each video segment corresponds to a total of 10 frames of task-state fMRI signals.
 Stimuli were presented using an LCD screen ($8^{\circ}\times 8^{\circ}$) positioned at the head of the scanner bed.
 Participants viewed the monitor via a mirror mounted on the RF coil and fixated a red central dot (0.4° × 0.4°).
-Preprocessing was performed using fMRIPrep ([10, 9]).
+Preprocessing was performed using fMRIPrep [12, 11].
 Following [36], the preprocessed functional data in 32k_fs_LR surface space were converted into 2D images and utilized for further analysis.
-Specifically, for each 32k_fs_LR data frame in each run, we applied z-scoring to the values across every vertex within the run.
-These z-scored values were then mapped onto 1023 × 2514 2D images using pycortex [11].
-The ROIs for analysis were selected from the HCP-MMP atlas [12] in the 32k_fs_LR space, including “V1, V2, V3, V3A, V3B, V3CD, V4, LO1, LO2, LO3, PIT, V4t, V6, V6A, V7, V8, PH, FFC, IP0, MT, MST, FST, VVC, VMV1, VMV2, VMV3, PHA1, PHA2, PHA3”.
+Given the delay of the BOLD signal by 6 seconds, we applied z-scoring to the data points across every vertex within each run, incorporating a 6-second lag.
+These normalized values were then projected onto 1023 × 2514 pixel 2D images using pycortex.
+For analysis, Regions of Interest (ROIs) were selected from the Human Connectome Project Multi-Modal Parcellation (HCP-MMP) atlas in the 32k_fs_LR space.
+These ROIs included areas such as “V1, V2, V3, V3A, V3B, V3CD, V4, LO1, LO2, LO3, PIT, V4t, V6, V6A, V7, V8, PH, FFC, IP0, MT, MST, FST, VVC, VMV1, VMV2, VMV3, PHA1, PHA2, PHA3”.
+Subsequently, the ROIs were converted into a 256 × 256 image, as illustrated in the right part of Fig.
+3.
 4 Problem Setup
-Figure 3: Overview of the MinD-3D Framework.
-Our approach combines a Neuro-Fusion Encoder for extracting features from fMRI frames, a Feature Bridge Diffusion Model for generating visual features from these fMRI signals, and a Latent Adapted Decoder based on the Argus 3D shape generator for reconstructing 3D objects.
-This integrated system effectively aligns and translates brain signals into accurate 3D visual representations.
-Note that the CLIP encoder is only for training the model, while it is not used for the inference stage.
 Recon3DMind addresses a task of paramount importance in cognitive neuroscience.
@@
 Consider the acquisition of a multi-frame fMRI signal denoted as $\{F\}$, where $|F|=10$.
-These signals correspond to both a 3D object point cloud, $\{P\}$, and a video, $\{V\}$, with $|V|=192$ frames, which the subject views.
+These signals correspond to both a 3D object mesh, $\Psi$, and a video, $\{V\}$, with $|V|=192$ frames, which the subject views.
 The task necessitates an efficient encoder, $E$, capable of extracting both spatial structural and semantic features from the fMRI signal.
@@
 Therefore, we input multiple frames of fMRI signals with the aim of extracting these comprehensive spatial structural features from the spatio-temporal data.
-This process is mathematically represented as: $f=E(F)$ Following the feature extraction, we employ a powerful decoder to reconstruct the original 3D point cloud $\{P\}$ based on the extracted feature $f$: $P=D(f)$ Consequently, our model can be succinctly defined as $M=\{E,D\}$, thereby operationalizing the transformation: $P=M(F)$.
+This process is mathematically represented as: $f=E(F)$ Following the feature extraction, we employ a powerful decoder to reconstruct the original 3D mesh $\Psi$ based on the extracted feature $f$: $\Psi=D(f)$ Consequently, our model can be succinctly defined as $M=\{E,D\}$, thereby operationalizing the transformation: $\Psi=M(F)$.
 In executing this model, it’s critical to fully utilize both the fMRI signal $\{F\}$ and the corresponding video $\{V\}$ to train our model $M$.
@@
 5 Method
+Figure 4: Overview of the MinD-3D Framework.
+Our approach combines a Neuro-Fusion Encoder for extracting features from fMRI frames, a Feature Bridge Diffusion Model for generating visual features from these fMRI signals, and a Latent Adapted Decoder based on the Argus 3D shape generator for reconstructing 3D objects.
+This integrated system effectively aligns and translates brain signals into accurate 3D visual representations.
+Note that the CLIP encoder is only for training the model, while not used for inference.
 5.1 Overview
 As depicted in Fig.
-3, our MinD-3D model comprises three main components: 1) Neuro-Fusion Encoder: This part is responsible for extracting both semantic and structural features from the fMRI frames.
+4, our MinD-3D model comprises three main components: 1) Neuro-Fusion Encoder: This part is responsible for extracting both semantic and structural features from the fMRI frames.
 2) Feature Bridge Diffusion Model: This module generates visual features based on the features extracted by the Encoder.
@@
 5.2 Neuro-Fusion Encoder
-Drawing from the potent Masked Image Model (MIM) [16, 26], we adopt an auto-encoder structure to extract meaningful fMRI features, trained on the UKB dataset [23].
-Our encoder, termed the Neuro-Fusion Encoder (NFE), utilizes the encoder of LEA [26] as $E_{f}$ with a Feature Aggregation module $\mathcal{FA}$ to transform fMRI frames into an embedding compatible with the CLIP [28] ViT-B/32 model’s space.
+Drawing from the potent Masked Image Model (MIM) [17, 13, 34], we adopt an auto-encoder structure to extract meaningful fMRI features, trained on the UKB dataset [30].
+Our encoder, termed the Neuro-Fusion Encoder (NFE), employs the encoder of LEA [34] as $E_{f}$, along with a Feature Aggregation module ($\mathcal{FA}$), to transform fMRI frames into embeddings.
+Inspired by contrastive learning [37, 9, 57], we then align these embeddings with the vision space through CLIP [37].
 Initially, we process each fMRI frame in parallel to obtain the spatial fMRI embeddings:
@@
 $\mathbf{c}_{v}=\frac{\sum^{n}_{i=1}E_{v}(\mathbf{V_{i}})}{n}$
-where n is the number of selected video frames. Then we utilize the CLIP loss to ensure the alignment between the fMRI features $\mathbf{c}_{f}$ and the visual features $\mathbf{c}_{v}$:
+where n is the number of selected video frames. Then we utilize the CLIP loss to ensure the alignment between fMRI features $\mathbf{c}_{f}$ and visual features $\mathbf{c}_{v}$:
 $\mathcal{L}_{c}=\mathcal{L}_{clip}(\mathbf{c}_{f},\mathbf{c}_{v})$
-During training, we optimize the whole Neuro-Fusion Encoder with $E_{LEA}$ loading the pre-trained weights and the CLIP vision encoder frozen.
+During training, we optimize the neuro-fusion encoder initialized with pre-trained weights, while the CLIP vision encoder remains frozen.
 It’s worth noting that the CLIP vision encoder $E_{v}$ and images $\{\mathbf{V}\}$ are just used for training, we will drop it during inference.
 By adopting this approach, we will get the fMRI features $\mathbf{c}_{f}$ and embedding $\mathbf{F}_{emb}$ aligned with the visual space.
-The fMRI features $\mathbf{c}_{f}$ will be used as the conditional information for the feature bridge diffusion model and fMRI embedding $\mathbf{F}_{emb}$ will be used in the latent adapted decoder.
-Figure 4: The qualitative results generated by LEA-3D, fMRI-PTE-3D, and our method are presented. GT indicates the ground-truth 3D objects. All the objects have been rendered into a 2D format.
-Methods Semantic-Level Structure-Level
-[…]
-LEA-3D[26] 0.787 0.371 0.353 0.527 0.562 0.575 0.624 0.372
-fMRI-PTE-3D[27] 0.815 0.392 0.375 0.433 0.694 0.327 0.519 0.135
-w/o Both 0.784 0.364 0.340 0.536 0.618 0.498 0.649 0.246
-w/o Diffusion 0.802 0.382 0.372 0.509 0.641 0.367 0.546 0.175
-w/o Contrastive 0.822 0.409 0.398 0.453 0.674 0.236 0.504 0.095
-MinD-3D (full) 0.835 0.425 0.411 0.419 0.704 0.214 0.480 0.065
-Table 1: Performance Comparison on the fMRI-Shape.
-We report average metrics for all subjects.
-LEA-3D and fMRI-PTE-3D represent variants of LEA and fMRI-PTE, respectively.
-The following 3 baselines are degenerate versions of MinD-3D.
+The fMRI features $\mathbf{c}_{f}$ will be used as the conditional information for the feature bridge diffusion model.
 5.3 Feature Bridge Diffusion Model
-This phrase details the Neural-Visual Synthesis process using a diffusion model, named here as the Feature Bridging Diffusion Model (FBDM), following the design outlined in [29].
-This module’s primary function is to facilitate the generation of the target visual latent feature $\hat{\mathbf{c}_{v}}$ from the corresponding fMRI latent feature $\mathbf{c}_{f}$.
+This section describes the Neural-Visual Synthesis process via a transformer-based diffusion model, herein referred to as the Feature Bridge Diffusion Model (FBDM), inspired by the hierarchical design proposed in [38].
+The FBDM’s primary role is to bridge fMRI latent features $\mathbf{c}_{f}$ with their visual counterparts $\hat{\mathbf{c}_{v}}$, effectively transforming neurological signals into visual representations.
 In the forward diffusion process, we treat the visual latent feature $\mathbf{c}_{v}$ as $\mathbf{x}_{0}$ and introduce Gaussian noise into it across 100 timesteps.
@@
 $\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{t-1}+\sqrt{1-\alpha_{t}}\mathbf{\epsilon},\quad\mathbf{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$
-where, $\mathbf{\epsilon}$ is the Gaussian noise.
-During the reverse diffusion process, the module will learn to predict and then reverse the noise introduced during the forward process.
-The goal is to accurately estimate the noise $\hat{\mathbf{\epsilon}}_{t}$ at each timestep, conditioned on the fMRI latent feature $\mathbf{c}_{f}$:
+During the reverse diffusion process, FBDM is tasked with predicting and reversing the introduced noise to reconstruct the latent visual feature.
+This is achieved by estimating the noise $\hat{\mathbf{\epsilon}}_{t}$ at each timestep, conditioned on the fMRI latent feature $\mathbf{c}_{f}$:
 $\hat{\mathbf{\epsilon}}_{t}=FBDM(\mathbf{x}_{t},t,\mathbf{c}_{f})$
-The fMRI latent feature $\mathbf{c}_{f}$ is mapped to the intermediate layers of the FBDM through a cross-attention operation:
-CrossAttention$(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)\cdot V$, with
-$Q=W_{Q}^{(i)}\varphi_{i}\left(x_{t}\right),K=W_{K}^{(i)}\tau_{\theta}(\mathbf{c}_{f}),V=W_{V}^{(i)}\tau_{\theta}(\mathbf{c}_{f}).$
-Here, $\tau_{\theta}$ is the fMRI latent feature with a suitable dimension projector, $\varphi_{i}\left(x_{t}\right)$ denotes intermediate values of the Decoder and $W_{Q}^{(i)}$, $W_{K}^{(i)},W_{V}^{(i)}$ are projector matrices with learnable parameters.
-So the training objective of FBDM is to minimize the difference between the actual and predicted noise:
+Differing from U-Net based models, our approach utilizes a transformer to leverage the self-attention mechanism for integrating the conditional information $\mathbf{c}_{f}$.
+This is done by embedding the timestep $t$, the noise-perturbed feature $\mathbf{x}_{t}$, and the conditional fMRI feature $\mathbf{c}_{f}$ into a unified representation, further augmented with a learnable token to capture the target vision features accurately.
+The objective of FBDM training is to minimize the discrepancy between the actual noise and its prediction:
 $\mathcal{L}_{FBDM}=\mathbb{E}_{\mathbf{x},\mathbf{\epsilon},t}\left[\|\mathbf{\epsilon}-\hat{\mathbf{\epsilon}}_{t}\|^{2}\right]$
-Post-training, the visual latent feature $\mathbf{c}_{v}$ will be dropped and FBDM initiates a generative phase, starting with pure noise and progressively applying the reverse diffusion process to produce the predicted visual latent feature $\hat{\mathbf{c}_{v}}$.
-This phase crucially maps the aligned fMRI latent features $\mathbf{c}_{f}$ to the visual space and generates the corresponding visual latent features $\hat{\mathbf{c}_{v}}$, which is used as the basic conditional information for the latent adapted decoder.
+Upon training completion, the model discards the initial visual latent feature $\mathbf{c}_{v}$ and enters a generative phase.
+It begins with pure noise and progressively refines this through the reverse diffusion process to generate the predicted visual latent feature $\hat{\mathbf{c}_{v}}$.
+This process effectively translates the aligned fMRI latent features $\mathbf{c}_{f}$ into visual representations, serving as essential conditional information for the subsequent latent-adapted decoding phase.
+Figure 5: The qualitative results generated by LEA-3D, fMRI-PTE-3D, and our method are presented. GT indicates the ground-truth 3D objects. All the objects have been rendered into a 2D format.
 5.4 Latent Adapted Decoder
-Latent Adapted Decoder is based on Argus [39], a cutting-edge 3D shape generator.
-It will generate the VQ latent representation of the target 3D object based on the conditional information we provide via a generative auto-regressive transformer and then get the 3D objects through the VQ decoder.
-Here, we design an adapter layer in the transformer block to modify the module using the fMRI embeddings $\mathbf{F}_{emb}$.
-Concretely, to directly incorporate the fMRI feature into the 3D generation process, we adapt the Multihead Self-Attention (MSA) in the standard transformer block into a double attention layer.
-This layer consists of the original MSA and an adapter layer.
-The adapter layer calculates multi-head cross-attention using the intermediate values of the module.
-We then obtain the adapted result using a learnable parameter $\lambda$ as the coefficient:
-$x=\lambda\cdot\textsc{CA}(Q,K_{f},V_{f})+(1-\lambda)\cdot\textsc{SA}(Q,K,V)$
-where CA and SA indicate cross-attention and self-attention separately; $Q,K,V$ denotes the representation of 3d latent features with a Linear projection; $K_{f},V_{f}$ are the representation of fMRI embeddings with a linear projection.
-The adapter layer is added every T transformer block.
-Our decoder computes the VQ latent vector for the target object:
-$p(\mathbf{z})=\prod_{i=1}^{m}p\left(\mathbf{z}_{i}\mid\hat{\mathbf{c}_{v}},\mathbf{F}_{emb},\mathbf{z}_{<i}\right)$
-Finally, the 3D VQ-decoder decodes this latent vector $\mathbf{z}$ to reconstruct the final point clouds $\mathbf{P}$:
-$\{\mathbf{P}\}=D(\mathbf{z})$
-The decoder’s optimization objective is defined by the negative log-likelihood loss:
+The Latent Adapted Decoder (LAD) leverages Argus [35], a state-of-the-art 3D mesh generation method, as its backbone.
+This component is tasked with generating the Vector Quantized (VQ) latent representation of target 3D objects from conditional information processed through a generative auto-regressive transformer.
+The resulting 3D objects are then reconstructed using the VQ decoder.
+To tailor Argus for processing fMRI input and to retain its generative efficiency, we employ adapter tuning during training.
+This involves integrating the adapter layers within the transformer block.
+Following classic methods [20, 53], we apply two feed-forward linear layers to effectively adapt the self-attention outputs.
+The adapter layer is systematically inserted at every $T$ transformer blocks, ensuring that the model remains adaptable and efficient.
+The decoder formulates the VQ latent vector for the target object as follows:
+$p(\mathbf{z})=\prod_{i=1}^{m}p\left(\mathbf{z}_{i}\mid\hat{\mathbf{c}_{v}},\mathbf{z}_{<i}\right)$
+Subsequently, the 3D VQ-decoder decodes this latent vector $\mathbf{z}$ into the final 3D mesh representation $\Psi$:
+$\Psi=D(\mathbf{z})$
+The decoder aims to minimize the negative log-likelihood loss during optimization:
 $\mathcal{L}_{nll}=\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}\left[-\log p(\mathbf{z})\right]$
-During training, we train the module separately using pre-training weights, after training the neuro-fusion encoder and the feature bridge diffusion model.
-6 Expriment
+For training, the decoder is fine-tuned from pre-trained weights, with the trained neuro-fusion encoder and the feature bridge diffusion model.
+This strategy preserves Argus’s generative capabilities while adapting to new inputs and minimizing the computational overhead associated with training.
+Table 2: Performance Comparison on fMRI-Shape.
+We report the average metrics for each subject trained and tested on their own data in the core part.
+LEA-3D and fMRI-PTE-3D represent variants of LEA and fMRI-PTE, respectively.
+The following three baselines are ablation studies for MinD-3D.
+Methods Semantic-Level Structure-Level
+[…]
+LEA-3D[34] 0.787 0.371 0.527 0.562 4.229 2.291 5.347
+fMRI-PTE-3D[36] 0.815 0.392 0.433 0.694 3.571 1.992 4.621
+w/o Both 0.789 0.367 0.479 0.616 3.694 2.205 5.073
+w/o Diffusion 0.801 0.385 0.423 0.669 3.526 2.071 4.625
+w/o Contrastive 0.823 0.419 0.319 0.701 3.315 1.826 4.027
+MinD-3D (full) 0.839 0.432 0.230 0.734 3.157 1.742 3.833
+6 Experiments
 6.1 Metrics
-In order to evaluate the performance of the new task and our model comprehensively, we adopt the following metrics in two aspects, semantic-level and structure-level:
+To more effectively evaluate the performance of our models on the new task, we utilize metrics in two aspects: semantic-level and structure-level.
 Semantic Level.
-To evaluate the semantic quality of our model, we adopt standard metrics used in previous 2D fMRI literature [24, 4, 5, 25, 22], namely pixel-level correlation (PixCorr) and N-way top-K accuracy.
-These two metrics are calculated by comparing the reconstructed images with the ground truth (GT) images.
-In this process, we render both the reconstructed objects and the GT objects into images separately, calculate the metrics for each frame, and then derive the average value as the final metric numbers.
+To evaluate the semantic quality of our model, we adopt standard metrics used in previous 2D fMRI literature [6, 7, 31, 29, 45], namely, N-way top-K accuracy.
+We calculate 2-way-top-1 and 10-way-top-1 accuracy as shown in Tab.
+2.
+Additionally, we calculate the Learned Perceptual Image Patch Similarity (LPIPS) [55] to assess the perceptual quality of semantics.
+In practice, these two metrics are calculated by comparing the reconstructed images with the ground truth (GT) images.
+In this process, we render both the reconstructed and GT objects into images separately at every 60 degrees, calculate the metrics for each frame, and then derive the average value as the final metric numbers.
 Structure Level.
-Besides evaluating the semantic quality of our model, it is also essential to assess its performance in terms of spatial structure, which we do in both 3D and 2D.
-In 3D, we utilize common metrics used in 3D reconstruction: Chamfer Distance (CD), Hausdorff Distance (HD), and Earth Mover’s Distance (EMD).
-For the 2D evaluation, we render the 3D objects onto a 2D plane and compute the frame-averaged pixel-level correlation, Learned Perceptual Image Patch Similarity (LPIPS), and Structural Similarity (SSIM).
-6.2 Implementation Details.
-As described in Section 3, each data pair in the fMRI-Shape dataset consists of 10 fMRI frames and the corresponding 192 images.
-To maximize the utilization of this dataset, we randomly select 6 fMRI frames for training, which also serves as a form of data augmentation.
-These 6 frames, initially 1023 × 2514 2D fMRI images, are resized to 256 × 256.
-For the image data, we set the number of selected video frames (n) to 4, meaning we randomly choose 4 frames for subsequent training, and each of these images is resized to 224 × 224.
-During inference, we select only the middle 6 fMRI frames and do not input any additional images.
-In all experiments, we train the model using data from all 3 subjects and then calculate the metrics for each subject.
-Regarding the model structure, our neuro-fusion encoder comprises 24 transformer layers with 1024 feature dimensions and 16 self-attention heads.
-Utilizing the CLIP ViT-B/32 model, our Feature Aggregation module maps features to a 512-dimensional space.
-Consequently, the feature bridge diffusion model is configured with a feature dimension of 512 for cross-attention and 8 self-attention heads, with the number of diffusion channels set to 224.
-The latent adapted decoder is designed with 32 transformer layers, 3072 feature dimensions, and 16 self-attention heads.
-We set T=8, which means that we add an adapter layer every 8 transformer blocks.
-In terms of model training, the first stage involves optimizing both the neuro-fusion encoder and the feature bridge diffusion model.
-We initialize the neuro-fusion encoder with pre-trained weights, and the feature bridge diffusion model is trained from scratch.
-In the second stage, we fix the parameters of the first two stages and focus on optimizing the latent adapted decoder.
-We use the AdamW optimizer with a learning rate of 5e-5 for all experiments and train the model with a batch size of 128 for all subjects.
+In addition to the semantic quality evaluation, it is crucial to assess the performance of our model in capturing geometrical structures.
+We utilize common metrics employed in 3D reconstruction [28, 52, 35]: Fréchet Point Cloud Distance (FPD, adjusted by a factor of ×$\times 10^{-1}$), Chamfer Distance (CD, scaled by ×$\times 10^{2}$), and Earth Mover’s Distance (EMD, also scaled by ×$\times 10^{2}$).
+We sample point clouds from the GT and reconstructed meshes to calculate these three metrics.
+Furthermore, we compute the frame-averaged Structural Similarity Index (SSIM) using the same rendering method to assess similarity from each viewpoint.
+6.2 Implementation Details
+As detailed in Sec.
+3, each data pair in the fMRI-Shape dataset is composed of 10 fMRI frames and 192 corresponding images.
+To maximize dataset utilization and introduce data augmentation, we randomly select 6 fMRI frames for training.
+These frames, originally 1023 × 2514 2D fMRI images, are resized to 256 × 256.
+For the image data, 8 frames are randomly selected for training, with each image resized to 224 × 224.
+During inference, only the middle 6 fMRI frames are used, without incorporating additional images.
+Regarding the architecture, we set $T=4$ to incorporate adapter layers.
+The model training proceeds in two stages.
+Initially, we focus on optimizing the neuro-fusion encoder and the feature bridge diffusion model.
+The neuro-fusion encoder is initialized with pre-trained weights, while the feature bridge diffusion model undergoes training from scratch, requiring approximately 2.5 days on a single A100 GPU.
+In the second stage, we optimize all parameters, including those of the adapters, and freeze the parameters in Argus.
+This stage is completed in about one day on a single A100 GPU.
+Details on the hyperparameters of the model are provided in the supplementary material.
 6.3 Experimental Results
-As the first work to model 3D imaging in the human brain, we compare our comprehensive model, MinD-3D, with several baselines, including adaptations of LEA and fMRI-PTE using the same 3D decoder as in MinD-3D.
+As the pioneering effort in modeling 3D imaging within the human brain, which involves various regions of the brain, direct comparisons with existing models are not feasible.
+To mitigate the training costs, we have adapted LEA and fMRI-PTE models, which were trained on the same vision ROIs, utilizing the identical 3D decoder as employed in MinD-3D.
+These adaptations serve as the baselines in this study, facilitating a more contextual and fair comparison within the novel domain we are exploring.
 Tab.
 2 presents the averaged metrics at both the Structural and Semantic Levels for all subjects.
-MinD-3D demonstrates superior performance across all metrics: achieving approximately 83.5% accuracy in 2-way-top-1, 42.5% accuracy in 10-way-top-1, a pixel-level correlation of 0.411, and 0.419 in LPIPS.
+MinD-3D demonstrates superior performance across all metrics: achieving approximately 83.9% accuracy in 2-way-top-1, 43.2% accuracy in 10-way-top-1, and scores of 0.734, 0.230, 3.157, 1.742, and 3.833 in other metrics.
 These results indicate that our model not only generates objects with high semantic accuracy but also excels in structural similarity, outperforming baseline methods.
+Tab.
+2 also proves the effectiveness of the diffusion model and the contrastive learning module, both of which play important roles in our MinD-3D.
 Qualitative results from LEA-3D, fMRI-PTE-3D, and our method are showcased in Fig.
-4.
+5.
 MinD-3D consistently produces 3D objects that are structurally very similar to their real counterparts, while maintaining semantic integrity in most instances.
@@
 This underscores the robustness of our model in handling this novel task and its capability to produce faithful realizations.
-6.4 Analytical and Ablation Study
-We further conduct some analytical and ablation studies to evaluate the effectiveness of our model.
-Figure 5: Quality of the extracted features.
-We map the features to brain regions via ridge regression modeling, with voxel-wise Pearson’s correlation coefficients computed between them.
-The visual function ROI is highlighted in red, indicating correlations in the range of 0 to 0.8.
-In contrast, other regions are depicted in blue, representing correlations within a range of 0 to 0.5.
-Analysis of feature extracted.
-To understand our feature extractor from a biological perspective, following [14, 37, 35, 24], we employ a linear encoding model here to project the extracted fMRI feature $c_{f}$ to brain activity.
-For this purpose, we select subject 1 as our representative.
-Specifically, we employ ridge regression to learn the mapping relationship between $c_{f}$ and fMRI signals in the training set.
-We then map $c_{f}$ extracted from the test set to the fMRI space and calculate the Pearson correlation coefficient.
-As illustrated in Fig.
-5, we use red to indicate the positive correlation between visual ROIs.
-For other regional positive correlations, we use blue.
-A high positive correlation is evident between $c_{f}$ and visual ROIs, indicating that we have successfully extracted valid features.
-Additionally, we conducted a similar analysis with the visual latent feature $c_{v}$ and visualized its positive correlations.
-Notably, the correlation of $c_{f}$ is superior to that of $c_{v}$.
-Furthermore, we calculated the correlation between $c_{f}$ and $c_{v}$, which is very high.
-This high correlation suggests that our comparative learning is effective and that the fMRI feature space and image space are well-aligned.
-Ablation study of fMRI frames.
-During training, we randomly select 6 fMRI frames, denoted as $F_{n}$, as the input to our model.
-To explore the impact of different $F_{n}$ values on our model’s performance, we conducted an assessment across various $F_{n}$ values.
-We report and compare average results for metrics such as 2-way-top-1, 10-way-top-1, SSIM, and CD, focusing on the performance across three subjects.
-As shown in Tab.
-2, $F_{n}=6$ emerges as the optimal choice.
-This operation not only acts as data augmentation but also, due to the inherent noise in fMRI signals, choosing the middle portion for inference helps in reducing the introduction of noise.
+6.4 Further Analysis
+To effectively utilize a subset of the fMRI-Shape dataset and assess the generalization capabilities of our proposed MinD-3D model, we conducted two Out-Of-Distribution (OOD) experiments under challenging settings: 1) Across-Person Testing (APT): In APT, we evaluate our model, which was trained only on Subject 1, using the data from Subject 9.
+We compare the results with the baselines and report the metrics in Tab.
+3.
+2) Across-Person & Across-Class Testing (APACT): In APACT, we similarly evaluate our model, trained solely on Subject 1, with the data from Subject 11.
+We also compare with the baselines and report the metrics in Tab.
+3.
+Although the performance in these OOD scenarios does not match the In-Distribution (ID) results, which is to be expected for such an extremely challenging task and given the substantial individual differences and domain gaps, our method still outperforms the baselines.
+This achievement not only underscores the robustness of MinD-3D but also establishes a new benchmark for the community to build upon.
 7 Conclusion
-$F_{n}$ Semantic-level Structure-Leve
+Table 3: Quantitative results of APT and APACT. We use the model trained on Subject 1 and compare metrics for Subjects 9 and 11 separately.
+Methods APT APACT
 […]
-1 0.674 0.291 0.472 0.853
-3 0.762 0.365 0.604 0.526
-6 0.835 0.425 0.704 0.214
-8 0.829 0.401 0.702 0.225
-10 0.818 0.379 0.684 0.273
-Table 2: Ablation results for the fMRI frames. It indicates that $F_{n}=6$ is the best choice for MinD-3D.
-In this paper, for the first time, we introduce the task of 3D reconstruction based on fMRI signals, accompanied by a dedicated dataset.
-Technically, we propose a novel three-stage framework tailored for this groundbreaking task.
-In the initial stage, our model adeptly extracts features from fMRI frames.
-Subsequently, in the second stage, it employs a diffusion module to translate these fMRI features into the image domain.
-The final stage involves the transformation of these image features into 3D models using an advanced 3D generation model.
-Our experimental results and analyses demonstrate the efficacy of our model in accurately extracting fMRI features and converting them into their corresponding 3D objects.
-This work not only establishes a new domain in neuroimaging and 3D reconstruction but also opens avenues for further research in understanding and visualizing neural representations in 3D vision.
-Appendix A More Implementation Details
+LEA-3D 5.362 3.627 6.174 6.958 4.944 8.107
+fMRI-PTE-3D 4.501 2.956 5.772 6.261 4.570 7.843
+MinD-3D 3.838 2.415 5.117 5.689 4.181 7.194
+In this paper, we introduce the innovative task of Recon3DMind and its accompanying large-scale dataset, fMRI-Shape, across various settings for the first time.
+Technologically, we develop a novel three-stage framework that involves more brain regions, including those associated with 3D vision in humans, specifically designed for this task.
+This approach sets a new benchmark in the field and proves the feasibility of the task.
+Initially, our model proficiently extracts features from fMRI frames.
+It then utilizes a diffusion module in the second stage to transition these fMRI features into the vision domain.
+The culmination of this process is the transformation of vision features into 3D models by an advanced 3D generation model in the final stage.
+Our comprehensive experimental results and analyses affirm the model’s effectiveness in precisely extracting fMRI features and converting them into their corresponding 3D objects.
+This pioneering work not only carves out a new niche in neuroimaging and 3D reconstruction but also paves the way for future research aimed at a deeper understanding and visualization of neural representations in 3D vision.
+Acknowledgements
+The computations in this research were performed using the CFFF platform of Fudan University.
+Appendix 0.A FMRI-Shape Details
+0.A.1 Visualization of fMRI Across Different Subjects
+To more effectively demonstrate brain activation patterns and showcase our fMRI-Shape dataset, we analyze and visualize the responses to five distinct objects across six different subjects in Fig.
+0.A.1.
+(Note: Only voxels with activation levels exceeding the 50th percentile are displayed.)
+Figure 0.A.1: Individual differences in brain activation patterns within the fMRI-Shape.
+Fig.
+0.A.1 also reveals significant individual differences in brain activation across subjects, more so than the variations in responses to different objects.
+This observation underscores the challenges inherent in the settings of AP and APAC.
+0.A.2 Visualization of the Core Component
+To illustrate the diversity of fMRI-Shape, Fig. 0.A.2 shows the core part, organized with five objects of training data and two objects of testing data each category.
+0.A.3 Visualization of the APAC Set
+The APAC Set, which includes additional objects not existing in the Core parts, is showcased in Fig. 0.A.3, with one object displayed for each category.
+Figure 0.A.2: Samples of Core component in fMRI-Shape. We show five objects of training data and two objects of testing data per category.
+Figure 0.A.3: Samples of APAC Set in fMRI-Shape. One object is shown per category.
+Appendix 0.B Analysis of Visualizations
+0.B.1 Visualizations from Different Subjects
+We display several samples from 5 subjects in Fig.
+0.B.1 to demonstrate the effectiveness of our MinD-3D model.
+Due to the varying responses of different subjects to different objects, there are some variations in the results.
+Overall, our model accurately reconstructs both the category and characteristics of the objects.
+The experiment validates the feasibility of our proposed task, Recon3DMind, and the accompanying dataset, fMRI-Shape, designed for this task.
+0.B.2 Visualizations of AP & APAC Testing
+We also showcase the reconstructed objects from AP & APAC testing in Fig.
+0.B.2.
+As highlighted in Fig.
+0.A.1, individual differences significantly affect the outcomes.
+Our AP & APAC Testing empirically confirms this.
+Despite the task’s high difficulty, MinD-3D successfully recovers the basic shapes of the objects, establishing a baseline method for the community.
+Figure 0.B.1: Samples from different subjects.
+Figure 0.B.2: Visualization of AP & APAC testing. AP Testing trains on Subject 1 and tests on Subject 9. APAC Testing trains on Subject 1 and tests on Subject 11.
+0.B.3 Fail Cases
+We present some failure cases in Fig. 0.B.3. Although the reconstruction does not always accurately capture the entire object, it still manages to retain the shape or characteristics of the objects.
+Figure 0.B.3: Fail cases.
+Appendix 0.C More Implementation Details
 To more effectively demonstrate the details of our model and its implementation, this section introduces the necessary preliminary knowledge and provides a detailed description of our neuro-fusion encoder and latent adapted decoder.
-A.1 Neuro-Fusion Encoder
+0.C.1 Neuro-Fusion Encoder
+Figure 0.C.1: Architecture of LEA. During training, LEA masks all the patch-based latent codes and enables the decoder to learn to reconstruct fMRI signals from the learned [CLS] token.
 Our encoder consists of the encoder from the pre-trained LEA [34] and a feature aggregation module.
 LEA, a transformer-based auto-encoder, encodes fMRI signals and reconstructs them as depicted in Fig.
-A.1.
-Unlike the original Masked Autoencoder (MAE) [16], which only masks some patch-based latent codes, LEA masks all such codes.
+0.C.1.
+Unlike the original Masked Autoencoder (MAE), which only masks some patch-based latent codes, LEA masks all such codes.
 This enables the decoder to learn to reconstruct fMRI signals from the learned [CLS] token during training, thereby facilitating the development of a dense fMRI representation within the [CLS] token.
@@
 As shown in Fig.
-A.1, when LEA is directly applied to our fMRI-Shape dataset, it yields rough reconstructions.
+0.C.1, when LEA is directly applied to our fMRI-Shape dataset, it yields rough reconstructions.
 However, these reconstructions are not highly precise, primarily due to variations in scanners, subjects, and tasks.
-Figure A.1: Architecture of LEA. During training, LEA masks all the patch-based latent codes and enables the decoder to learn to reconstruct fMRI signals from the learned [CLS] token.
 Given that our input fMRI signals are in video form, to efficiently extract both semantic and structural features from this complex data, we opt to fine-tune the encoder with a small learning rate during the initial stage of training with hyperparameters provided in Tab.
-A.1.
+0.C.1.
+Table 0.C.1: Hyperparameters used in neuro-fusion encoder.
 parameter value parameter value
@@
 mlp ratio 0.99 image frames 4
-Table A.1: Hyperparameters used in neuro-fusion encoder.
-A.2 Latent Adapted Decoder
-Our decoder is adapted from Argus [39].
+0.C.2 Feature Bridge Diffusion Model
+We use a transformer-based diffusion model to generate vision features.
+During the inference phase of the diffusion model, we concatenate the fMRI token, the time token, the predicted image token, and the learnable token.
+The output learnable token then serves as the predicted image token for the model’s next inference.
+The hyperparameters of our feature bridge diffusion model are shown in Tab.
+0.C.2.
+Table 0.C.2: Hyperparameters used in feature bridge diffusion model.
+parameter value parameter value
+depth 6 heads 8
+embedding dim 512 timestamps 100
+mlp ratio 0.99 token length 770
+0.C.3 Latent Adapted Decoder
+Our decoder is adapted from Argus.
 The original Argus model enhances the quality and diversity of generated 3D shapes by incorporating tri-plane features as latent representations and a discrete codebook for efficient quantization, leveraging the power of transformers for multi-modal conditional generation.
-It generates latent codes in an auto-regressive manner, similar to VQGAN [8], based on conditional features.
-Subsequently, it decodes the quantized vectors into point clouds.
-In our latent adapted decoder, we add an adapter layer after every 8 transformer blocks to fuse the fMRI feature.
-During the second stage of training, we focus on optimizing the parameters in the adapter layers and the learnable parameter $\lambda$.
+It generates latent codes in an auto-regressive manner, similar to VQGAN, based on conditional features.
+Subsequently, it decodes the quantized vectors into 3D mesh.
+In our latent-adapted decoder, we add an adapter layer after every two (T=2) Transformer blocks to fuse the fMRI feature.
+During the second stage of training, we freeze the parameters of Argus and focus on optimizing the parameters in the adapter layers.
 The hyperparameters of our latent adapted decoder are detailed in Tab.
-A.2.
+0.C.3.
+Table 0.C.3: Hyperparameters used in latent adapted decoder.
 parameter value parameter value
@@
 embedding dim 3072 adapter layers 4
-mlp ratio 0.99 sequence length 1024
-Table A.2: Hyperparameters used in latent adapted decoder.
-Appendix B FMRI-Shape Details
-To better illustrate our proposed dataset, fMRI-Shape, we provide a detailed description in this section.
-Figure B.1: Examples of 360-Degree View Videos.
-These videos effectively showcase the intricate details, lighting, and shadows of the 3D objects.
-Presented in a sequence from left to right, they depict the order in which the subjects viewed them.
-Each video boasts a high resolution of 1024×1024, providing a clear and detailed visual experience.
-B.1 More Experimental Details
-To clarify the experimental details, we showcase our specially rendered 360-degree view videos.
-In Fig.
-B.1, six frames are selected from four videos.
-The camera in each video encircles the object, completing a full 360-degree clockwise rotation, forming an 8-second video with 192 frames.
-These high-resolution videos (1024×1024 pixels) vividly capture the intricate details, lighting effects, and shadows of the 3D objects.
-The arrangement of the videos mirrors the order in which they were viewed by the subjects.
-Displayed on an LCD screen, these videos enable the subjects to observe and comprehend all facets of the 3D objects.
-This viewing strategy is crucial, as it aids the subjects’ brains in accurately perceiving the objects, a fundamental aspect for gathering precise and high-quality fMRI signals that correspond to the visual stimuli observed by the subjects.
-B.2 More Preprocessing Details
-In preparing our data for training, we provide a visualization that clearly shows how we identify the visual Regions of Interest (ROIs) in our data. The preprocessing is applied to all fMRI frames.
-Figure B.2: Selection Process of Visual ROIs: On the left, the fMRI is depicted in its original signal form, while on the right, the corresponding area is transformed into a 2D image, demonstrating the conversion process.
-Refer to Fig.
-B.2.
-The first row displays the raw fMRI signals and their 2D image formation.
-In the second row, on the left, you can see the appearance of the fMRI signal and the locations of the ROIs.
-On the right, there is a 2D image illustrating where these ROIs are located within the signal, which has been converted from a format known as 32k_fs_LR.
-After identifying these ROIs in the 1023×2514 2D fMRI data, we resize them to 256×256, as shown in Fig.
-B.3.
-It’s important to note that our Mind-3D model uses these resized, single-channel 2D fMRI signals as its input.
-Figure B.3: Converting ROIs in fMRI Signal
-Appendix C More Results
-Subject Semantic-Level Structure-Level
-[…]
-sub-001 0.842 0.424 0.411 0.421 0.704 0.212 0.478 0.069
-sub-002 0.836 0.424 0.412 0.420 0.705 0.218 0.483 0.062
-sub-003 0.829 0.427 0.410 0.416 0.703 0.212 0.479 0.064
-Table C.1: Quantitative results for all subjects.
-Qualitative results from different views.
-In Fig.
-C.1, we showcase a collection of 2D rendered images of 3D objects by MinD-3D from various viewpoints for subject-1 in fMRI-Shape, closely resembling the ground truth (GT) objects both in semantics and structure.
-Quantitative results for all subjects.
-Our proposed dataset, fmri-shape, includes data from three subjects.
-To demonstrate the robustness of our model, we present comprehensive quantification results for all three subjects.
-Tab.
-C.1 showcases these full metrics, indicating that our model consistently performs well across different subjects.
-These results underscore the model’s robustness in handling subject variability.
-Figure C.1: Qualitative results for subject-1 from various viewpoints.
-Appendix D Limitations and Future Works
-As the first attempt at the task of Reconstructing 3D Objects from Mind in fMRI (Recon3DMind), our MinD-3D has successfully reconstructed 3D objects that are semantically and structurally similar to the originals.
-However, the reconstructed 3d objects lack detailed texture, color, and other aspects compared to the original point clouds.
-To overcome this limitation, our future work will focus on developing methods to more effectively extract texture and color details from 3D objects within fMRI signals.
-We aim to refine our reconstruction method to create more accurate and detailed replications of the original 3D objects, thereby enhancing the fidelity of the reconstructed models and capturing the complexity and richness of the original designs.
-Regarding our dataset, we recognize that relying solely on the fMRI-Shape dataset is insufficient for the task’s full scope.
-A notable challenge is mitigating dataset noise due to the cognitive biases individuals might hold towards certain objects.
-Therefore, it’s necessary for us to expand our fMRI-Shape dataset by including more 3D objects and collecting data from additional subjects, with the hope of advancing research in our proposed Recon3DMind field.
+mlp ratio 0.99 sequence length 1027
+Appendix 0.D Features Analysis
+To understand the biological relevance of our feature extractor, in line with previous studies [51, 47, 32], we employ a linear encoding model to project the extracted fMRI features $c_{f}$ onto brain activity patterns, using subject 1 as a representative case for analysis.
+Figure 0.D.1: Quality of the extracted features.
+We map the features to brain regions via ridge regression modeling using $c_{f}$, $c_{v}$ and $\hat{c_{v}}$, with voxel-wise Pearson’s correlation coefficients computed between them.
+The visual function ROI is highlighted in red.
+In contrast, other regions are depicted in blue.
+Specifically, we utilize ridge regression to establish the mapping between $c_{f}$ and the fMRI signals within the training dataset.
+We then project $c_{f}$ from the test set into the fMRI space and compute the Pearson correlation coefficient to assess the relationship.
+As illustrated in Fig.
+0.D.1, areas positively correlated with visual Regions of Interest (ROIs) are marked in red, while other regions showing positive correlations are highlighted in blue.
+The results demonstrate a strong positive correlation between $c_{f}$ and visual ROIs, confirming the effective extraction of relevant features.
+Additionally, we perform a similar correlation analysis with the visual latent feature $c_{v}$, which also shows positive correlations.
+Notably, the correlation strength of $c_{f}$ surpasses that of $c_{v}$, indicating its superior relevance to brain activity.
+However, it is important to note that in some measures, the correlation coefficient of $c_{v}$ is higher than that of $c_{f}$, which highlights the complexity of the task and underscores the critical role of bridging features.
+Furthermore, our results indicate that our generated visual latent features achieve activation levels comparable to the ground truth (GT), emphasizing the effectiveness of our comparative learning approach and demonstrating well-aligned fMRI feature and image spaces.
+Appendix 0.E Future Works
+As the first attempt at the task of Reconstructing 3D Objects from the Mind using fMRI (Recon3DMind).
+Our proposed MinD-3D model has successfully reconstructed 3D objects that are semantically and structurally similar to their originals, thereby proving the feasibility of the task.
+However, our model’s 3D representations, which are mesh-based, accurately outline the shape of objects but do not capture color and other details as precisely as the original objects.
+To improve on this, we plan to develop new methods for more effectively extracting texture and color information from fMRI signals.
+Our goal is to refine our reconstruction technique to produce more detailed and accurate reproductions of the original 3D objects, thereby improving the quality of the models and more faithfully capturing the originals’ complexity and nuances.
+Crucially, we introduce the first 3D vision fMRI dataset with contributions from 14 participants across various settings.
+Currently, most fMRI-based visual reconstruction methods are specific to individuals, making out-of-distribution (OOD) testing, both across individuals and classes, extremely challenging.
+Our proposed fMRI-Shape dataset presents an opportunity to tackle this formidable challenge.
+Addressing individual differences is a complex task that necessitates a collaborative effort from the research community with the goal of enhancing the adaptability of such models.
+Regarding our dataset, we acknowledge that relying solely on the fMRI-Shape dataset is not sufficient for comprehensively addressing the task at hand.
+Therefore, we plan to enrich the dataset by sampling more 3D data, including all objects from ShapeNet, and involving more participants.
+This effort aims to expand the fMRI-Shape dataset and further advance research in fMRI-based human brain decoding.